OpenStudy (anonymous):

write the equation of the line which passes through (2,-1) and is perpendicular to the line with equation 2y-x=1

OpenStudy (anonymous):

and how would i do this ?

OpenStudy (anonymous):

1º. Put in slope intersect form 2y-x=1------1/2x +1 =y 2º slope of perpendicular line is-2, so 3º its equation Y+1 = -2(X-2)

OpenStudy (anonymous):

you have ... y=k(x-x_1) you have... 2y-x=1 2y=1+x \[\Large y=\frac12 +\frac x2\] \[\Large y=\boxed{\frac 12}x+\frac12 \] the selected one is k now you have the formula \[\LARGE k\cdot k_1=-1\] when is perpendicular and \[\LARGE k=k_1\] when is parallel but the second one we don't need now we have.. \[\LARGE k\cdot k_1=-1\] \[\LARGE k\cdot \frac12=-1\] \[\LARGE k=-2\] now we have the formula... \[\LARGE (y-y_1)=k(x-x_1)\] substitute and you'll get... \[\LARGE (y-(-1))=-2(x-2)\] \[\LARGE (y+1))=-2x+4\] go on...

OpenStudy (anonymous):

no idea ...

OpenStudy (anonymous):

none of the answer choices look like that and i dont know if you mean go on like there is another step

hero (hero):

@aanders please go back to your previous question regarding absolute value. I posted a complete solution for you. Next time, don't leave your question so quickly and always get more than one opinion. Just because someone says your answer is correct, that doesn't mean it is correct.

OpenStudy (anonymous):

good advice thank you